(x+2, 2y-3)=(0,1) so (x,y)=?
Answer by jim_thompson5910(35256)

(x+2, 2y-3)=(0,1) ---> x+2 = 0 and 2y-3 = 1

x+2 = 0

x = -2

-----------

2y -3 = 1

2y = 1+3

2y = 4

y = 4/2

y = 2

So (x,y) = (-2, 2)